A particle moves with a velocity 6i-4j+3k m/s under the influence of a constant force F=20i+15j-5k B. The influence power applied to the particle is?

Grade:11

1 Answers

hamza
6 years ago
using formulaepower = force ×velocityP=F×vP=(20i+15j-5k)×(6i-4j+3k)now multiplyingP=(120i-6j-15k)